OpenStudy (anonymous):

I really need help figuring out how to do this i can't keep relying on open study lol plz help Find the x-coordinate of each point where the graph of the function f(x)=x^2(2x+3)^2 has a horizontal tangent line

OpenStudy (accessdenied):

You know how to find the derivative of the function?

OpenStudy (anonymous):

ya

OpenStudy (accessdenied):

Okay, that would be the first step. The derivative is the slope of the tangent. For what slope would we have a horizontal line?

OpenStudy (anonymous):

y=something

OpenStudy (anonymous):

do i use product rule and chain rule ?

OpenStudy (accessdenied):

to find the derivative, yes. and yeah, that would be the equation itself... the slope of that horizontal line is 0.

OpenStudy (anonymous):

so i need to let u=2x+3 and f(u)=2x^2(2x+3)^2

OpenStudy (anonymous):

oops f(u) should be 2x^2(u)^2

OpenStudy (anonymous):

and i then just apply product rule?

OpenStudy (anonymous):

so here is what I've done with chain rule u=2x+3 f(u)=2x^2(u)^2 so, dy/dx=dy/du(du/dx) ((4x)(2(u)))(2) ((4x)(4x+6))(2) (16x^2+24x)(2) 32x^2+48x

OpenStudy (anonymous):

now do i just plug that into product rule with the rest of the equation?

OpenStudy (accessdenied):

we'd use product rule first... since this would be the product of two functions, f(x) = x^2 and g(x) = (2x + 3)^2 Derivative of f(x)*g(x) = f'(x) g(x) + f(x) g'(x). We need chain rule only for the derivative of g(x)

OpenStudy (anonymous):

oh ok, that makes since i feel retorted i should have known that.

OpenStudy (accessdenied):

its fine, we learn from our mistakes. :)

OpenStudy (anonymous):

then (x^2)d/dx(2x_3)^2+(2x+3)d/dx(x^2)

OpenStudy (anonymous):

oops it should be (x^2)d/dx(2x_3)^2+(2x+3)^2d/dx(x^2)

OpenStudy (anonymous):

and i only use chain rule to find out the derivative of (2x+3)^2 right

OpenStudy (accessdenied):

right.

OpenStudy (anonymous):

k thanks I'm gonna see how all that works out for me

OpenStudy (accessdenied):

or if you really wanted, you could multiply out your f(x) and then take the derivative only using power rule, since we're going to simplify it into a polynomial form anyways.

OpenStudy (accessdenied):

that way would actually be easier, i just didnt realize it. :P

OpenStudy (anonymous):

oh well I've actually already done it using chain rule.. but i came up with (x^2)(8x+12)^2+(2x)(2x+3)^2

OpenStudy (accessdenied):

(8x+12)^2 i dont agree with the g(x) = (2x + 3)^2, with the outer function being x^2 and the inner being 2x+3...

OpenStudy (anonymous):

well i got the 8x+12 using chain rule haha

OpenStudy (anonymous):

u=2x+3 h(u)=u^2 =(2u)(2) =(2(2x+3))(2) =(4x+6)2 =8x+12

OpenStudy (accessdenied):

okay, that parts correct. you just squared your answer in the equation, which didn't make sense...

OpenStudy (anonymous):

ya because its squared in the original equations x^2(2x+3)^2

OpenStudy (accessdenied):

it is squared there, but the derivative of that part is not. 8x + 12 is the exact derivative of g(x), so back in the product rule, you'd insert it exactly as such, the derivative of g(x).

OpenStudy (anonymous):

ok so it is (x^2)(8x+12)+(2x)(2x+3)

OpenStudy (accessdenied):

(x^2)(d/dx((2x+3)^2)) + d/dx( x^2 )(2x + 3)^2 = (x^2)(8x+12)+(2x)(2x+3)^2 ^----------------equals----------------^ the (2x + 3) would be squared still, since that's the original equation I believe you were thinking of.

OpenStudy (anonymous):

ya ok that makes sense so only the term that we took the derivative of loses its square

OpenStudy (accessdenied):

essentially, yes

OpenStudy (anonymous):

ok so now that we have the product rule set out once i solve for that how do i get a final numerical answer

OpenStudy (accessdenied):

well, since we now have the derivative function, we'll go back to that "horizontal tangent" thing. We want to find the points where the derivative, or slope of the tangent, is 0. So, we can just set this up by setting the derivative equal to 0 and solving for x. (x^2)(8x+12)+(2x)(2x+3)^2 = 0 We'll need to simplify the left-hand side first, getting it into standard polynomial form. So, we'd multiply our the (2x + 3)^2, etc.

OpenStudy (accessdenied):

2x(2x + 3)^2 + (x^2)(8x + 12) = 2x(4x^2 + 12x + 9) + (8x^3 + 12x^2) = 8x^3 + 24x^2 + 18x + 8x^3 + 12x^2 = 16x^3 + 36x^2 + 18x that's what I get

OpenStudy (accessdenied):

Now, if we set this equal to 0... 16x^3 + 36x^2 + 18x = 0 we can factor out a monomial (2x) 2x(8x^2 + 18x + 9) = 0 then divide off the 2, we need factors of 8*9 that add up to 18. x(8x^2 + 18x + 9) = 0 72 factors into 1+72, 2+36, 3+24, 4+18, 6+12, and 8+9. 6+12 adds up to 18, so we'll split the 18x up and factor by grouping. x(8x^2 + 6x + 12x + 9) = 0 x(2x(4x + 3) + 3(4x + 3)) = 0 x(4x + 3)(2x + 3) = 0 Then 0-product property, a*b = 0; a=0, b=0 to solve. x = 0 4x + 3 = 0 4x = -3 x = -3/4 2x + 3 = 0 2x = -3 x = -3/2 And we're done! x = 0, -3/4, and -3/2 are the x-coordinates of the points where we have horizontal tangents.

OpenStudy (anonymous):

im not sure why but thats actually what i got and al three of those answers are wrong.. at least according to my online homework..

OpenStudy (anonymous):

nevermind man i had to add the two x coordinates together thats the right answer